A lonely guy throws a stone into a still pond causing a circular ripple to spread. If the radius
of the circular ripple spreads at the rate 1.5 ft/sec, how fast is the enclosed area increasing at the end of 2 seconds?
1
Expert's answer
2020-12-09T08:26:31-0500

Area of circle =πr2..... A = πr2


Differentiating it with respect to time..

"\LARGE\frac{dA}{dt}" =2πr "\LARGE\frac{dr}{dt}"

If the radius is increasing at a constant rate of "\LARGE\frac{1.5 ft}{sec}"

then after 2 seconds radius = 3 fts.


We know that "\LARGE\frac{dr}{dt}" = "\LARGE\frac{1.5ft}{sec}"


and so "\LARGE\frac{dA}{dt}" = [2]π[3][2]

"=\LARGE 12π\,ft^2/sec" .